OpenStudy (anonymous):

If the first equation is multiplied by 3 and then the equations are added, the result is? 3x + y = 3 x - 3y = -2

OpenStudy (anonymous):

where are you confused? If you wanna share your answer for the first part (multiplying by 3) i can help check your work.

OpenStudy (anonymous):

i got 10x=11

OpenStudy (anonymous):

Not quite... Here let me help out by writing it how you should do the math. \[\huge 3(3x+y=3) \equiv 3(3x+y) = 3*3\]

OpenStudy (anonymous):

9x + y = 9 9x + y=9 ??

OpenStudy (anonymous):

9xy 9xy=9

OpenStudy (anonymous):

You first answer is close! you're missing something with 'y'

OpenStudy (anonymous):

Oh and the 3 lines is just a sign that means "equivalent" i was just trying to show two steps, not have two separate problems. does that make sense?

OpenStudy (anonymous):

yea is it 10x = 7

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

thanks for the help!